# Skew 2-Dyck paths via the kernel method

### v1

@@ Abstract.
We continue on a recent concept introduced by Kariuki and Okoth, about skew 2-Dyck paths, introducing an additional down-step $L$, together with the usual steps $U$ (up) and $D$ down. There is the syntactical condition that ULUL and LULU can never occur. An automaton that checks these conditions is introduced, and the relevant generating functions are obtained by applying the kernel method to three functional equations. It is briefly discussed how the setting can be extended to $t$-Dyck paths. As a benefit, prefixes of skew $t$-Dyck paths are also enumerated.
@@
The new paper by Kariuki and Okoth [6] offers the new idea to introduce left down-steps. If L=(1,2)L=(-1,-2) this is not very interesting, as, related to the first paper on skew Dyck paths by Deutsch et al. [2] one should be careful that the path does not intersect (overlaps) itself, which in such a setting it wouldn’t do. Thus it is the right idea to stretch down-steps to D=(2,2)D=(2,-2) and L=(2,2)L=(-2,-2). Note that down-steps resp. down-left-steps still only count as one step. As we did before in our generating function based analysis of skew Dyck paths [3], instead of (2,2)(-2,-2) we draw (2,2)(2,-2) in red color. The syntactic conditions to be fulfilled are that ULUL and LULU are forbidden. A step $L$ can only follow a $D$-step, and cannot be followed by an $U$-step; note that the very last step may be an $L$-step.
@@
The rest of this paper is devoted to an analysis (=enumeration) of skew 2-Dyck paths, using the kernel method. Some recent applications of these techniques can be found in [4], [5], [3]. There are some advantages: The analysis is in a sense straight forward; one does not need to invent clever decompositions of the paths in question, and one gets the enumeration of partial skew 2-Dyck paths as a bonus, as we have have all three generating functions leading to level $i$ by an up-step, down-step, red down-step in an explicit form.
In a last section, we indicate how to extend the concept from skew 2-Dyck paths to skew $t$-Dyck paths, again handling prefixes as a bonus.
@@ 2. Automaton to syntactically check 2-Dyck paths with left (red) steps
We start by constructing an automaton that checks whether the syntacting conditions are still fulfilled, i. e., ULUL and LULU are forbidden. According to the three types of steps, we have three layers of states (current level).
@@
For the enumeration, we introduce $f_{i}=f_{i}(z)$, $g_{i}=g_{i}(z)$, $h_{i}=h_{i}(z)$, the generating functions of skew 2-Dyck paths leading to state $i$ in the first (second, third) layer, where the exponent $z^{k}$ counts the number of steps when starting at the origin (thick node on the first layer). The recursions are as follows:
@@
They are easy to understand by distinguishing the last possible step that leads to level ii. Further, one introduces bivariate generating functions
@@
These 4 solutions where found using gfun and its algeqtoseries construction. [^1] Thanks to Cyril Banderier for valuable hints.
@@
$h_{0}$ $=\frac{1}{zs_{4}}-\frac{z}{s_{4}^{2}}-1={z}^{6}+6{z}^{9}+34{z}^{12}+198{z}^{15}+1191{z}^{18}+7364{z}^{21}+\dots,$
@@
Surprisingly, the numbers of skew 2-Dyck paths of length $3n$ have a simple generating function (seq A007564):
@@
The extraction of coefficient is easy when using the substitution $z=\dfrac{t}{(1+t)(1+3t)}$; then $R=1+t$, and contour integration (or the Lagrange inversion formula) leads to
@@ 3. Generalization: skew $t$-ary Dyck-paths
We start with skew 3-Dyck paths, from which the general case about skew $t$-Dyck paths will become clear almost immediately. There is the usual up-step $U$, and a down-step $D$ of 3 units, and a red down-step $L$ of 3 units; the conditions are that ULUL and LULU are forbidden.

### v2

@@ Abstract.
We continue on a recent concept introduced by Kariuki and Okoth, about skew 2-Dyck paths, introducing an additional down-step $L$, together with the usual steps $U$ (up) and $D$ (down). There is the syntactical condition that $UL$ and $LU$ can never occur. An automaton that checks these conditions is introduced, and the relevant generating functions are obtained by applying the kernel method to three functional equations. It is briefly discussed how the setting can be extended to $t$-Dyck paths. As a benefit, prefixes of skew $t$-Dyck paths are also enumerated. An approach that scans 2-Dyck paths from right to left is also discussed.
@@
The new paper by Kariuki and Okoth [7] offers the new idea to introduce left down-steps. If L=(1,2)L=(-1,-2), this is not very interesting, as, related to the first paper on skew Dyck paths by Deutsch et al. [2] one should be careful that the path does not intersect (overlaps) itself, which in such a setting it wouldn t do. Thus it is the right idea to stretch down-steps to D=(2,2)D=(2,-2) and L=(2,2)L=(-2,-2). Note that down-steps resp. down-left-steps still only count as one step. As we did before in our generating function based analysis of skew Dyck paths [3], instead of (2,2)(-2,-2) we draw (2,2)(2,-2) in red color. The syntactic conditions to be fulfilled are that $UL$ and $LU$ are forbidden. A step $L$ can only follow a $D$-step, and cannot be followed by an $U$-step; note that the very last step may be an $L$-step.
@@
The rest of this paper is devoted to an analysis (=enumeration) of skew 2-Dyck paths, using the kernel method. Some recent applications of these techniques can be found in [4], [5], [3]. There are some advantages: The analysis is in a sense straight forward; one does not need to invent clever decompositions of the paths in question, and one gets the enumeration of partial skew 2-Dyck paths as a bonus, as we have all three generating functions leading to level $i$ by an up-step, down-step, red down-step in an explicit form.
In a further section, we indicate how to extend the concept from skew 2-Dyck paths to skew $t$-Dyck paths, again handling prefixes as a bonus.
In a last section, 2-Dycks are scanned from right to left. If they return to $x$-axis, the enumeration is of course the same as before.
@@ 2. Automaton to syntactically check 2-Dyck paths with left (red) steps
We start by constructing an automaton that checks whether the syntactic conditions are still fulfilled, i. e., $UL$ and $LU$ are forbidden. According to the three types of steps, we have three layers of states (current level).
@@
For the enumeration, we introduce $f_{i}=f_{i}(z)$, $g_{i}=g_{i}(z)$, $h_{i}=h_{i}(z)$, the generating functions of skew 2-Dyck paths leading to state $i$ in the first (second, third layer, where the exponent $z^{k}$ counts the number of steps when starting at the origin (thick node on the first layer). The recursions are as follows:
@@
They are easy to understand by distinguishing the last possible step that leads to level $i$. Further, one introduces bivariate generating functions
@@
These solutions where found using gfun and its algeqtoseries construction. [^1] Thanks to Cyril Banderier for valuable hints.
@@
$h_{0}$ $=\frac{1}{zs_{4}}-\frac{z}{s_{4}^{2}}-1={z}^{6}+6{z}^{9}+34{z}^{12}+198{z}^{15}+1191{z}^{18}+7364{z}^{21}+\dots\;.$
In total
@@
Surprisingly, the numbers of skew 2-Dyck paths of length $3n$ have a simple generating function (seq A007564 in [6]):
@@
The extraction of coefficients is easy when using the substitution $z=\dfrac{t}{(1+t)(1+3t)}$; then $R=1+t$, and contour integration (or the Lagrange inversion formula) leads to
@@ 3. Generalization: skew $t$-ary Dyck-paths
We start with skew 3-Dyck paths, from which the general case about skew $t$-Dyck paths will become clear almost immediately. There is the usual up-step $U$, and a down-step $D$ of 3 units, and a red down-step $L$ of 3 units; the conditions are that $UL$ and $LU$ are forbidden.
@@
4. 2-Dyck paths read from right to left
First, we mirror the previous Figure 3. For the enumeration of 2-Dyck paths this mirrored version must lead to the same results. However, when considering partial 2-Dyck paths, results will be different.
Figure 6. All 2-Dyck paths of 9 steps, with possible red down steps indicated. Right-left version.
The previous automaton can be used, but all arrows must be reversed.
01122334455667788
Figure 7. Graph (automaton) to recognize skew 2-Dyck paths. The first step of a non-empty path must be an up-step or a red up-step. Start can be in the second or third layer; the empty path can only be counted once.
We set up the usual recursions, use the same letters, but now with a different meaning.
$f_{0}$ $=0,\quad f_{1}=zf_{2},\quad f_{i+2}=zg_{i}+zf_{i+3},\ i\geq 0,$
$g_{0}$ $=1+zf_{1},\quad g_{1}=zf_{2},\quad g_{i+2}=zf_{i+3}+zg_{i}+zh_{i},\ i\geq 0,$
$h_{0}$ $=1,\quad h_{1}=0,\quad h_{i+2}=zg_{i}+zh_{i},\ i\geq 0.$
Summing,
$\sum_{i\geq 0}u^{i+3}f_{i+2}$ $=\sum_{i\geq 0}u^{i+3}zg_{i}+\sum_{i\geq 0}u^{i+3}zf_{i+3},$
$\sum_{i\geq 0}u^{i+3}g_{i+2}$ $=\sum_{i\geq 0}u^{i+3}zf_{i+3}+\sum_{i\geq 0}u^{i+3}zg_{i}+\sum_{i\geq 0}u^{i+3}zh_{i},$
$\sum_{i\geq 0}u^{i+2}h_{i+2}$ $=\sum_{i\geq 0}u^{i+2}zg_{i}+\sum_{i\geq 0}u^{i+2}zh_{i},$
or
$u(F-uf_{1})$ $=u^{3}zG+z(F-uf_{1}-u^{2}f_{2}),$
$u(G-g_{0}-ug_{1})$ $=z(F-uf_{1}-u^{2}f_{2})+u^{3}zG+u^{3}zH,$
$H-1$ $=u^{2}zG+u^{2}zH.$
The solution is
$F$ $=\frac{u(2zu^{2}g_{0}-zu^{2}-g_{0}+1)}{-2{u}^{3}z+{z}^{2}{u}^{2}+{z}^{3}{u}^{4}+u-z},$
$G$ $=\frac{-z^{2}u^{2}+u-zg_{0}+z^{2}u^{2}g_{0}}{-2{u}^{3}z+{z}^{2}{u}^{2}+{z}^{3}{u}^{4}+u-z},$
$H$ $=\frac{u-z-z^{2}u^{2}g_{0}}{-2{u}^{3}z+{z}^{2}{u}^{2}+{z}^{3}{u}^{4}+u-z}.$
The factorization of the denominator $2u^{3}z+z^{2}u^{2}+z^{3}u^{4}+u-z$ is easy since it is the reciprocal denominator from before, where $u\leftrightarrow u^{-1}$. Thus we have
$2u^{3}z+z^{2}u^{2}+z^{3}u^{4}+u-z=z^{3}(u-s_{1}^{-1})(u-s_{2}^{-1})(u-s_{3}^{-1})(u-s_{4}^{-1}).$
The only ‘bad’ factor is $(u-t_{1})$ with $t_{1}=s_{1}^{-1}$. Dividing it out from the numerator of $F$, we get $g_{0}+1+2t_{1}^{2}zg_{0}-t_{1}^{2}z$, but since $f_{0}=0$, we find finally
$g_{0}=\frac{1-zt_{1}^{2}}{1-2zt_{1}^{2}}=-1-zt_{1}^{2}+\frac{2t_{1}}{z}=1+z^{3}+4z^{6}+19z^{9}+100z^{12}+563z^{15}+3322z^{18}+20285z^{21}+\dots\!.$
The expressions $u^{k}]G(u)$ can be computed as well, but are not as pleasant as before since we have to deal with a factor $\dfrac{1}{(u-s_{2}^{-1})(u-s_{3}^{-1})(u-s_{4}^{-1})}$, and extraction of $u^{k}$ is not nice, although doable when applying partial fraction decomposition first. We leave details to interested readers.